OpenStudy (anonymous):

Solve the system using the elimination method 3x+2y=11 and 4x+y=-2

OpenStudy (anonymous):

so what we want to do is get one of the variables to cancel out

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

im thinking the easiest way would be to multiply the second equation by 2

OpenStudy (anonymous):

-2 sorry

OpenStudy (anonymous):

in my opinion the elimination methood is so much easier than the substitution method lol

OpenStudy (anonymous):

depends on the problem can go either way

OpenStudy (anonymous):

so do you understand how to do this?

OpenStudy (anonymous):

ok so i X -2 and go -8x+y=-4 is that correct?

OpenStudy (anonymous):

you need to multiply everything by -2 so you would get -8x-2y=4

OpenStudy (anonymous):

that would be a -4 correct because -2x2 = -4? :)

OpenStudy (anonymous):

-2*-2=4

OpenStudy (anonymous):

-2 (4x+y=-2)=-8x-2y=4

OpenStudy (anonymous):

im sorry caculator error my bad you're right it equals 4. now that i have that what is my next step

OpenStudy (anonymous):

add the two equations together

OpenStudy (anonymous):

the y's should cancel out and leave you with x. simplify and you have your x value

OpenStudy (anonymous):

i added them together and got -5x and is it -5x + or - 4y=7?

OpenStudy (anonymous):

3x+2y=11 +-8x-2y=4 -5x=15

OpenStudy (anonymous):

so what does x=?

OpenStudy (anonymous):

what answer do you get before the step -5x=15?

OpenStudy (anonymous):

what step are you talking about?

OpenStudy (anonymous):

after you add then together what do you do

OpenStudy (anonymous):

what's the step before -5x=15?

OpenStudy (anonymous):

\[(3x+2y=11)+(-8x-2y=4)=-5x=15\]

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yes thank you so idk if i may ask this or not but do you happen to have a facebook account? lol

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

may i look you up or no?

OpenStudy (anonymous):

if you look up my name you should be able to find it, same picture

OpenStudy (anonymous):

ya thats fine

OpenStudy (anonymous):

okay what's your name

OpenStudy (anonymous):

Nick Relich

OpenStudy (anonymous):

cool you live in green bay W?

OpenStudy (anonymous):

ya